# Biblio.bib

### v2

@@ Abstract.
We show that, given a real or complex hyperbolic metric $g_{0}$ on a closed manifold $M$ of dimension $n\geq 3$, there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the space of negatively curved metrics such that for any $g\in\mathcal{U}$, the topological entropy and Liouville entropy of $g$ coincide if and only if $g$ and $g_{0}$ are homothetic. This provides a partial answer to Katok’s entropy rigidity conjecture. As a direct consequence of our theorem, we obtain a local rigidity result for the hyperbolic rank and for metrics with $C^{2}$ Anosov foliations near complex hyperbolic metrics.
@@ 1.1. Statement of the problem
The metric $g$ lifts to a metric $g_{\mathrm{Sas}}$ on $SM$ called the Sasaki metric (see [Pat, Chapter 1]) and the norms in the previous inequalities are taken with respect to $g_{\mathrm{Sas}}$. The Sasaki metric defines a natural smooth Riemannian volume form $\mu_{\mathrm{Liou}}$ which is also the Liouville form associated to the contact structure of $SM$. The (normalized) measure $\mu_{\mathrm{Liou}}$ is called the Liouville measure, it is smooth and invariant by the flow.
An Anosov flow has infinitely many invariant probability measures. The following result is classical and known as the variational principle:
@@
Here, $\mathrm{Ent}_{\mathrm{top}}(\varphi_{1}^{g})$ denotes the topological entropy of the time one map and $\mathrm{Ent}(\varphi_{1}^{g},\mu)$ denotes the metric entropy of the time one map with respect to the measure $\mu$. Moreover, for an Anosov flow, the supremum is attained for a unique invariant measure $\mu_{\mathrm{BM}}$ for which we have $\mathrm{Ent}(\varphi_{1}^{g},\mu_{\mathrm{BM}})=\mathrm{Ent}_{\mathrm{top}}(g)$. This measure $\mu_{\mathrm{BM}}$ is the measure of maximal entropy (or Bowen-Margulis measure). A natural question, first raised by Katok in [Ka, Section 2], is to characterize the negatively curved metrics $g$ for which the Liouville measure and the measure of maximal entropy coincide. Note that because of the variational principle recalled above, this is equivalent to $\mathrm{Ent}_{\mathrm{Liou}}(g):=\mathrm{Ent}(\varphi_{1}^{g},\mu_{\mathrm{Liou}})=\mathrm{Ent}_{\mathrm{top}}(g)$. The following conjecture is known as Katok’s entropy rigidity conjecture [Ka, BK].
@@
Let $(M^{n},g_{0})$ be a closed manifold of dimension $n\geq 3$ and suppose that $g_{0}$ is either a real hyperbolic or a complex hyperbolic metric. Then there exists $N(n)\in\mathbb{N}$ and $\varepsilon>0$ such that for any negatively curved metric $g$ on $M$ with $\|g-g_{0}\|_{C^{N}}<\varepsilon$, if one has $\mathrm{Ent}_{\mathrm{Liou}}(g)=\mathrm{Ent}_{\mathrm{top}}(g)$ then $g$ is homothetic to $g_{0}$ and thus locally symmetric.
@@ 1.2. Existing results
Katok showed in [Ka] the conjecture on surfaces (i.e $n=2$) and his proof relied on the fact that any negatively curved metric is conformally equivalent to a hyperbolic metric in this case. His argument extends to any dimension for metrics in the conformal class of a locally symmetric metric but this does not exhaust all metrics in dimension $n\geq 3$.
We note that Foulon [Fou] and then De Simoi, Leguil, Vinhage and Yang [SLVY] generalized the result of Katok. In their setting, the geodesic flow on the 3-dimensional manifold $SM$ is replaced by a general Anosov flow on a 3-manifold. The Liouville measure is replaced by the contact volume or a smooth invariant measure and they show that equality with the measure of maximal entropy implies that the system is conjugated to an algebraic flow (see [Fou, SLVY] for the precise statements).
In higher dimensions, fewer results are known and there has been no progress on this conjecture for more than two decades. Let us first state a major result of Besson, Courtois, Gallot [BCG] which asserts that for a closed manifold $M^{n}$ which admits a locally symmetric metric $g_{0}$ (unique in this case by Mostow’s rigidity theorem), then $g\mapsto\mathrm{Ent}_{\mathrm{top}}(g)$ is (strictly) minimized at $g_{0}$ among negatively curved metrics $g$ with volume equal to that of $g_{0}$. For the entropy of the Liouville measure, Flaminio [Fla, Theorem C] showed that there exist hyperbolic metrics in dimension $3$ which neither maximize nor minimize $g\mapsto\mathrm{Ent}_{\mathrm{Liou}}(g)$ among metrics of constant volume, see also [Mau] for a generalization to higher dimensions.
Flaminio in [Fla, Theorem A] obtained an ”infinitesimal” version of Katok’s conjecture near real hyperbolic metrics in any dimension. To state his result, let us first introduce a functional $\Phi$ defined on negatively curved metrics:
@@
In particular, the functional $\Phi$ is constant along any orbit $\mathcal{O}(g)$ and the kernel of the Hessian $d^{2}\Phi(g_{0})$ always contains the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. If $g_{0}$ is locally-symmetric, then $\Phi(g_{0})=0$ and $d\Phi(g_{0})=0$ as $g_{0}$ is a critical point of $\Phi$. Then Flaminio proved that for a real hyperbolic metric in any dimension $n\geq 3$, the Hessian $d^{2}\Phi(g_{0})$ is a positive quadratic form with kernel given exactly by $T_{g_{0}}\mathcal{O}(g_{0})$. This infinitesimal version of Katok’s entropy conjecture shows that if we consider a smooth deformation $(g_{\lambda})_{\lambda\in(-\varepsilon,\varepsilon)}$ of $g_{0}$ (in the space of negatively curved metrics of same volume) for which one has $\Phi(g_{\lambda})\equiv 0$, then the deformation is tangent to the orbit of $g_{0}$ at $g_{0}$. See also [PaPe, Corollary 2.3] for a local result near metrics in the conformal class of a real hyperbolic metric (with an assumption on the curvature).
@@ 1.3. Entropy rigidity near locally symmetric metrics
In the following paper, the analysis is made ”transversally” to the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. For a nearby metric $g$, we will first need to ”project” it on $\mathrm{Ker}(D_{g_{0}}^{*}):=(T_{g_{0}}\mathcal{O}(g_{0}))^{\perp}$ (the space of divergence-free symmetric two-tensors, see Lemma 2.2). For $g$ close enough to $g_{0}$, its orbit $\mathcal{O}(g)$ intersects $\mathrm{Ker}(D_{g_{0}}^{*})$ at one single point which we will denote by $\phi_{g}^{*}g$, see Lemma 2.3 for a precise statement. We will denote by $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ the orthogonal projection from the space of symmetric tensors onto $\mathrm{Ker}(D_{g_{0}}^{*})$. Using microlocal methods, we prove the following result, which is valid for any locally symmetric metric.
@@
Let $(M^{n},g_{0})$ be a closed manifold with $n\geq 3$ and $g_{0}$ a locally symmetric negatively curved metric. Then there exists a differential operator $Q$ ((((made explicit in Theorem 6)))) and an open neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3n/2+7}$-topology, such that if $g\in\mathcal{U}$ satisfies $\phi_{g}^{*}g-g_{0}\in\mathrm{Ker}(\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})})^{\perp}$, one has
@@
The operator $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ is injective if $g_{0}$ is a real hyperbolic metric or if $g_{0}$ is a complex hyperbolic metric.
@@ 1.4. Local hyperbolic rank rigidity
A Riemannian manifold $(M,g)$ has higher hyperbolic rank if for any geodesic $\gamma$, there is a non-vanishing Jacobi field $J(t)$ along $\gamma$ such that $J(t)$ and $\dot{\gamma}(t)$ span a plane of curvature equal to $-1$. Higher hyperbolic rank on closed manifold is conjectured to only occur for locally symmetric spaces. We list thereafter some known results (see [CNS20] for a more detailed account of the existing literature):
  • if $(M,g)$ has higher hyperbolic rank and its sectional curvature satisfies $K\leq-1$ then it is a locally symmetric space of rank $1$ [Ham],
  • if $(M,g)$ has higher hyperbolic rank and its sectional curvature is 1/4-pinched (that is $-1\leq K\leq-1/4$) then it is a locally symmetric space of rank $1$ [CNS20, Theorem 1.1]. See also [Cons, Corollary 1] for similar results under different pinching conditions.
In [CNS20], the authors made the following conjecture.
@@
In a subsequent paper [CNS21], they prove the following local results near locally symmetric metrics.
  • Let $(M,g_{0})$ be either a closed quaternionic or octonionic hyperbolic locally symmetric manifold. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3}$-topology such that if $g\in\mathcal{U}$ has higher hyperbolic rank with sectional curvature satisfying $K\geq-1$, then it is a locally symmetric space [CNS21, Theorem 1.14].
  • Let $(M,g_{0})$ be a closed complex hyperbolic locally symmetric manifold. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3}$-topology such that if $g\in\mathcal{U}$ has higher hyperbolic rank with sectional curvature satisfying $K\geq-1$, then its Liouville entropy and topological entropy coincide [CNS21, Theorem 1.15].
@@
Theorem 4 (Local rigidity of hyperbolic rank near complex hyperbolic metrics).
@@
This is an analogue of [CNS21, Theorem 1.14] for complex hyperbolic metrics. Note however that with our techniques, the neighborhood is taken in the $C^{3n/2+7}$-topology while their results hold for $C^{3}$ metrics.
@@
The regularity of the Anosov splitting associated to the geodesic flow on a negatively curved manifold is a property that characterize locally-symmetric metrics along all negatively curved metrics. More precisely:
  • Hurder and Katok [HuKa] showed that for negatively curved surfaces, if the Anosov splitting is $C^{2}$, then it is $C^{\infty}$.
  • Kanai [Kan] showed that in dimension $n\geq 3$ and for $4/9$-pinched negatively curved metrics, smoothness of the Anosov splitting implied that $g$ is isometric to a real hyperbolic metric. See also Feres and Katok [FeKa1, FeKa2] for the same result with no pinching condition in dimension $3$ and the optimal pinching in dimension $4$.
  • Benoist, Foulon and Labourie [BFL] showed that in any dimension $n$, there exists $k(n)\in\mathbb{N}$ such that if a negatively curved metric has $C^{k}$ foliations, then it is locally symmetric.
The value $k$ in the previous theorem grows with the dimension $n$ and one can take $k(n)=2(2n^{2}-n+1)$. However, as noticed by Hamenstädt in [Ham2, Theorem B], for a negatively curved metric, if the Anosov splitting is $C^{2}$, then the Liouville measure is equal to the measure of maximal entropy. From Theorem 1, we thus obtain:
@@
Flaminio’s argument consists in using the geometry of real hyperbolic spaces to first simplify the expression of the Hessian $d^{2}\Phi(g_{0})$. We start by generalizing the computation to general locally symmetric metrics (see Theorem 3.1). Then Flaminio uses the representation theory of these manifolds to obtain the positivity and non-degeneracy of the Hessian outside of $T_{g_{0}}\mathcal{O}(g_{0})$.
The representation theory is replaced here by analytical techniques. More precisely, we use the ”generalized X-ray transform” whose ellipticity on suitable spaces of symmetric tensors was the core of the arguments of [GuLef, GuKnLef]. The basic properties of this operator are recalled in Subsection 2.2. Using this formalism seems to have two advantages over the representation theory argument:
  • it generalizes to other types of locally symmetric metrics (not only real hyperbolic ones),
  • we obtain a stability estimate (1.3) which shows a genuine injectivity result (not only along smooth deformations).
Acknowledgements. The author would like to first thank Colin Guillarmou and Thibault Lefeuvre for their careful and precious guidance in during the writing of this paper.
@@
Finally, the author would like to thank Andrei Moroianu for answering some questions on complex and quaternionic geometry.

### v3

# Katok’s entropy conjecture near real and complex hyperbolic metrics

@@ Abstract.
We show that, given a real or complex hyperbolic metric $g_{0}$ on a closed manifold $M$ of dimension $n\geq 3$, there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the space of negatively curved metrics such that for any $g\in\mathcal{U}$, the topological entropy and Liouville entropy of $g$ coincide if and only if $g$ and $g_{0}$ are homothetic. This provides a partial answer to Katok’s entropy rigidity conjecture. As a direct consequence of our theorem, we obtain a local rigidity result of the hyperbolic rank near complex hyperbolic metrics.
@@ 1.1. Statement of the problem
The metric $g$ lifts to a metric $g_{\mathrm{Sas}}$ on $SM$ called the Sasaki metric (see [31, Chapter 1]) and the norms in the previous inequalities are taken with respect to $g_{\mathrm{Sas}}$. The Sasaki metric defines a natural smooth Riemannian volume form $\mu_{\mathrm{Liou}}$ which is also the Liouville form associated to the contact structure of $SM$. The (normalized) measure $\mu_{\mathrm{Liou}}$ is called the Liouville measure, it is smooth and invariant by the flow.
An Anosov flow has infinitely many invariant probability measures. The following result is classical and known as the variational principle, see for instance [12, Corollary 4.3.9],
@@
Here, $\mathrm{Ent}_{\mathrm{top}}(\varphi_{1}^{g})$ denotes the topological entropy of the time-one map and $\mathrm{Ent}(\varphi_{1}^{g},\mu)$ denotes the metric entropy of the time one map with respect to the measure $\mu$. Moreover, for an Anosov flow, the supremum is attained for a unique invariant measure $\mu_{\mathrm{BM}}$ for which we have $\mathrm{Ent}(\varphi_{1}^{g},\mu_{\mathrm{BM}})=\mathrm{Ent}_{\mathrm{top}}(g)$. This measure $\mu_{\mathrm{BM}}$ is the measure of maximal entropy (or Bowen-Margulis measure). A natural question, first raised by Katok in [24, Section 2], is to characterize the negatively curved metrics $g$ for which the Liouville measure and the measure of maximal entropy coincide. Note that because of the variational principle recalled above, this is equivalent to $\mathrm{Ent}_{\mathrm{Liou}}(g):=\mathrm{Ent}(\varphi_{1}^{g},\mu_{\mathrm{Liou}})=\mathrm{Ent}_{\mathrm{top}}(g)$. The following conjecture is known as Katok’s entropy rigidity conjecture [24, 4].
@@
A negatively curved locally symmetric manifold is covered by a either a real, complex, quarternionic or octonionic hyperbolic space, see Subsection 2.3 for more details.
@@
Let $(M^{n},g_{0})$ be a closed manifold of dimension $n\geq 3$ and suppose that $g_{0}$ is either a real hyperbolic or a complex hyperbolic metric. Then there exists $N(n)\in\mathbb{N}$ and $\epsilon>0$ such that for any negatively curved metric $g$ on $M$ with $\|g-g_{0}\|_{C^{N}}<\epsilon$, if one has $\mathrm{Ent}_{\mathrm{Liou}}(g)=\mathrm{Ent}_{\mathrm{top}}(g),$ then $g$ is homothetic to $g_{0}$ and thus locally symmetric.
@@ 1.2. Existing results
Katok showed Conjecture 1 for surfaces (i.e., $n=2$) [24], and his proof relied on the fact that any negatively curved metric is conformally equivalent to a hyperbolic metric in this case. His argument extends to any dimension for metrics in the conformal class of a locally symmetric metric but this does not exhaust all metrics in dimension $n\geq 3$.
We note that Foulon [14], and then De Simoi, Leguil, Vinhage and Yang [35], generalized the result of Katok. In their setting, the geodesic flow on the 3-dimensional manifold $SM$ is replaced by a general Anosov flow on a 3-manifold. The Liouville measure is replaced by the contact volume or a smooth invariant measure, and they show that equality with the measure of maximal entropy implies that the system is conjugate to an algebraic flow (see [14, 35] for the precise statements).
In higher dimensions, fewer results are known. Let us first state a major result of Besson, Courtois, Gallot [2] which asserts that for a closed manifold $M^{n}$ which admits a locally symmetric metric $g_{0}$ (unique in this case by Mostow’s rigidity theorem), then $g\mapsto\mathrm{Ent}_{\mathrm{top}}(g)$ is (strictly) minimized at $g_{0}$ among negatively curved metrics $g$ with volume equal to that of $g_{0}$. For the entropy of the Liouville measure, Flaminio [13, Theorem C] showed that there exist hyperbolic metrics in dimension $3$ which neither maximize nor minimize $g\mapsto\mathrm{Ent}_{\mathrm{Liou}}(g)$ among metrics of constant volume, see also [29] for a generalization to higher dimensions.
Flaminio in [13, Theorem A] obtained an “infinitesimal” version of Katok’s conjecture near real hyperbolic metrics in any dimension. To state his result, let us first introduce a functional $\Phi$ defined on negatively curved metrics:
@@
In particular, the functional $\Phi$ is constant along any orbit $\mathcal{O}(g),$ and the kernel of the Hessian $d^{2}\Phi(g_{0})$ always contains the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. If $g_{0}$ is locally-symmetric, then $\Phi(g_{0})=0$ and $d\Phi(g_{0})=0$ as $g_{0}$ is a critical point of $\Phi$. Then Flaminio proved that for a real hyperbolic metric in any dimension $n\geq 3$, the Hessian $d^{2}\Phi(g_{0})$ is a positive quadratic form with kernel given exactly by $T_{g_{0}}\mathcal{O}(g_{0})$. This infinitesimal version of Katok’s entropy conjecture shows that if we consider a smooth deformation $(g_{\lambda})_{\lambda\in(-\epsilon,\epsilon)}$ of $g_{0}$ (in the space of negatively curved metrics of same volume) for which one has $\Phi(g_{\lambda})\equiv 0$, then the deformation is tangent to the orbit of $g_{0}$ at $g_{0}$.
@@ 1.3. Entropy rigidity near locally symmetric metrics
In this paper, the analysis is made “transversally” to the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. For a nearby metric $g$, we will first need to “project” it on $\mathrm{Ker}(D_{g_{0}}^{*}):=(T_{g_{0}}\mathcal{O}(g_{0}))^{\perp}$ (the space of divergence-free symmetric two-tensors, see Lemma 2.2). For $g$ close enough to $g_{0}$, its orbit $\mathcal{O}(g)$ intersects $\mathrm{Ker}(D_{g_{0}}^{*})$ at one single point which we will denote by $\phi_{g}^{*}g$, see Lemma 2.3 for a precise statement. We will denote by $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ the orthogonal projection from the space of symmetric tensors onto $\mathrm{Ker}(D_{g_{0}}^{*})$. Using microlocal methods, we prove the following result, which is valid for any locally symmetric metric.
@@
Let $(M^{n},g_{0})$ be a closed manifold with $n\geq 3$ and $g_{0}$ a locally symmetric negatively curved metric. Then there exists a differential operator $Q$ ((made explicit in Theorem 5)) and an open neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3n/2+7}$-topology, such that if $g\in\mathcal{U}$ satisfies $\phi_{g}^{*}g-g_{0}\in\mathrm{Ker}(\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})})^{\perp}$, one has
@@
The operator $\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}Q\Pi_{\mathrm{Ker}(D_{g_{0}}^{*})}$ is injective if $g_{0}$ is a real or complex hyperbolic metric.
@@ 1.4. Local hyperbolic rank rigidity
A Riemannian manifold $(M,g)$ has higher hyperbolic rank if for any geodesic $\gamma$, there is a non-vanishing Jacobi field $J(t)$ along $\gamma$ such that $J(t)$ and $\dot{\gamma}(t)$ span a plane of curvature equal to 1-1. Higher hyperbolic rank on closed manifold is conjectured to only occur for locally symmetric spaces. We list thereafter some known results (see [6] for a more detailed account of the existing literature):
  • if $(M,g)$ has higher hyperbolic rank and its sectional curvature satisfies $K\leq-1$ then it is a locally symmetric space of rank $1$ [20],
  • if $(M,g)$ has higher hyperbolic rank and its sectional curvature is 1/4-pinched (that is $1\leq K\leq-1/4$) then it is a locally symmetric space of rank 11 [6, Theorem 1.1]. See also [7, Corollary 1] for similar results under different pinching conditions.
In [6], the authors made the following conjecture.
@@
In a subsequent paper [5], they prove the following local results near locally symmetric metrics.
  • Let $(M,g_{0})$ be either a closed quaternionic or octonionic hyperbolic locally symmetric manifold. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3}$-topology such that if $g\in\mathcal{U}$ has higher hyperbolic rank with sectional curvature satisfying $K\geq-1$, then it is a locally symmetric space [5, Theorem 1.14].
  • Let $(M,g_{0})$ be a closed complex hyperbolic locally symmetric manifold. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{3}$-topology such that if $g\in\mathcal{U}$ has higher hyperbolic rank with sectional curvature satisfying $K\geq-1$, then its Liouville entropy and topological entropy coincide [5, Theorem 1.15].
@@
Corollary 1 (Local rigidity of hyperbolic rank near complex hyperbolic metrics).
@@
This is an analogue of [5, Theorem 1.14] for complex hyperbolic metrics. Note however that with our techniques, the neighborhood is taken in the $C^{3n/2+7}$-topology while their results hold for $C^{3}$ metrics.
@@
The regularity of the Anosov splitting associated to the geodesic flow on a negatively curved manifold is a property that characterizes locally-symmetric metrics along all negatively curved metrics. More precisely:
  • Hurder and Katok [22] showed that for negatively curved surfaces, if the Anosov splitting is $C^{2}$, then it is $C^{\infty}$.
  • Kanai [23] showed that in dimension $n\geq 3$ and for $4/9$-pinched negatively curved metrics, smoothness of the Anosov splitting implied that $g$ is isometric to a real hyperbolic metric. See also Feres and Katok [11, 10] for the same result with no pinching condition in dimension $3$ and the optimal pinching in dimension $4$.
  • Benoist, Foulon and Labourie [1] showed that in any dimension $n$, there exists $k(n)\in\mathbb{N}$ such that if a negatively curved metric has $C^{k}$ foliations, then it is locally symmetric.
The value $k$ in the previous theorem grows with the dimension $n$ and one can take $k(n)=2(2n^{2}-n+1)$. However, as noticed by Hamenstädt in [21, Theorem B], for a negatively curved metric, if the Anosov splitting is $C^{2}$, then the Liouville measure is equal to the measure of maximal entropy. From Theorem 1, we thus obtain:
@@
The first part of Flaminio’s argument consists in using the geometry of real hyperbolic spaces to first simplify the expression of the Hessian $d^{2}\Phi(g_{0})$. We start by generalizing the computation to all locally symmetric metrics (see Theorem 3.1). The computation relies on the geometrical characterization of locally symmetric spaces recalled in Subsection 2.3.
The main difference with Flaminio’s strategy is that we rewrite the Hessian using the “generalized X-ray transform” $\Pi$. The basic properties of the operator $\Pi$ are recalled in Subsection 2.2.
Note that the ellipticity and injectivity of $\Pi$ on suitable spaces of symmetric tensors was the core of the arguments of [17, 16]. Our proof share similarities with the ones of [17, 16]. In particular, the stability estimate (1.3) is deduced from a Taylor expansion once one shows the Hessian $d^{2}\Phi(g_{0})$ is elliptic and injective when restricted to solenoidal tensors, which we refer to as the infinitesimal problem. The ellipticity of the Hessian is deduced from the works of Guillarmou and Guillarmou-Lefeuvre [15, 17, Theorem 3.5, Lemma 4.3]) and an explicit computation of the principal symbol of the differential operator $Q$ appearing in Theorem 3.1, see Proposition 3.3.
The infinitesimal problem for real hyperbolic metrics was solved by Flaminio using a Weitzenböck formula. In Section 4.1, we show Theorem 3, that is, we solve the infinitesimal problem for real and complex hyperbolic metrics.
The proof of Theorem 3 for complex hyperbolic metrics is much more involved than for real hyperbolic metrics. The technical reason is that the differential operator $Q$ appearing in the Hessian computation takes values in symmetric $4$-tensors for the complex hyperbolic case while it takes values in symmetric $2$-tensors for the real hyperbolic one. Note that the infinitesimal problem corresponding to the marked length spectrum rigidity conjecture studied in [17, 16] was solved by Croke and Sharafutdinov [8] using ideas from tensor tomography and inverse problems. We will follow this approach to prove Theorem 3.
More precisely, we will use a correspondance between symmetric tensors and spherical harmonics recalled in Subsection 2.1. The complex structure allows to further decompose the spaces of spherical harmonics. We develop in Subsection 2.4 a theory of raising/lowering operators for general complex hyperbolic spaces very similar to the work of Guillemin–Kazhdan on surfaces [19] which might be of independent interest. In Section 4.1, we use this theory as well as a Weitzenböck formula to complete the proof of Theorem 3.
Acknowledgements. The author would like to first thank Colin Guillarmou and Thibault Lefeuvre for their careful and precious guidance during the writing of this paper.
@@
The author would like to thank Andrei Moroianu for answering some questions on complex and quaternionic geometry.
Finally, the author would like to thank the anonymous referee for their careful and thorough reading of the paper.
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (grant agreement No. 101162990).